OpenStudy (anonymous):

can some one help me with 6 rational expressions?

OpenStudy (anonymous):

\[\frac{3}{ z-4 }+\frac{6 }{ z+3 }\]

OpenStudy (anonymous):

what?

hero (hero):

@Spartan_Of_Ares, I could have sworn I showed you how to simplify rational expressions.

OpenStudy (anonymous):

omggg help me please

OpenStudy (anonymous):

i wanted to know if i did this right

OpenStudy (anonymous):

spartan

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats the thing is

hero (hero):

@jevans17, post your own question

OpenStudy (anonymous):

first i got \[\frac{ 3 }{ z-4 }\]\[\frac{ z+18 }{ (z+3)^2 }\] when multiplying

OpenStudy (anonymous):

that came out wrong and jevans i have no clue what your talking about

hero (hero):

Spartan, where's your teacher?

OpenStudy (anonymous):

no my problem is not an equation

OpenStudy (anonymous):

my teacher was sick today and we didn't really cover this

hero (hero):

Well, you should wait until tomorrow and ask her.

OpenStudy (anonymous):

\[\frac{ 3 }{ z-4 }+\frac{ z+18 }{ (z+3)^2 }\] this is how it came out i dont have her again until next tuesday

hero (hero):

That's not what you're supposed to do.

OpenStudy (anonymous):

ask your own questio jevans

hero (hero):

You're supposed to do this: \[\frac{3}{z-4} \times 1 + \frac{6}{z + 3} \times 1\] \[\left(\frac{3}{z-4} \times \frac{z+3}{z+3}\right) + \left(\frac{6}{z+3} \times \frac{z-4}{z-4}\right)\]

OpenStudy (anonymous):

heroo help me man

OpenStudy (anonymous):

help me

hero (hero):

@jevans17, where's your Geometry teacher?

OpenStudy (anonymous):

i have none i really need help

hero (hero):

I can't help fifteen people all at once.

OpenStudy (anonymous):

pelase help me rea quick sir please

hero (hero):

I still don't understand why you refuse to post your own question

OpenStudy (anonymous):

thats my question sir its thee same

OpenStudy (anonymous):

\[i got \frac{ 3z+9 }{ z^2-12 }+\frac{ 6z-24 }{ z^2-12 }\]

OpenStudy (anonymous):

:{

OpenStudy (anonymous):

lol he dumb

hero (hero):

@jevans17, go to your own question

OpenStudy (anonymous):

hero it is my own question its the same question

hero (hero):

@Spartan_Of_Ares, you got the numerators correct.

OpenStudy (anonymous):

on a scale of 1-10 how badly did i do that?

OpenStudy (anonymous):

real bacl a 10

OpenStudy (anonymous):

F

hero (hero):

The denominators need work. See if you can understand this: \[(z - 4)(z + 3) = z(z + 3) - 4(z + 3) = z^2 + 3z - 4z - 12 = z^2 - z - 12\]

OpenStudy (anonymous):

Hero help me pelase i dont have all day

OpenStudy (anonymous):

GTFO!!

hero (hero):

\[\space\space\space\space\space(z - 4)(z + 3) \\= z(z + 3) - 4(z + 3) \\= z^2 + 3z - 4z - 12 \\= z^2 - z - 12\]

OpenStudy (anonymous):

Ares shut the f up and learn to multiply

hero (hero):

@Spartan_Of_Ares, do you understand what I am posting?

OpenStudy (anonymous):

yes i just didnt understand how you were getting a trinomial but that make sense

OpenStudy (anonymous):

Hero how do i get someone else who can help me cause your no help at all

OpenStudy (anonymous):

post your own question!

hero (hero):

@jevans17, I refuse to help someone who infiltrates someone else's question.

OpenStudy (anonymous):

I posted my own

OpenStudy (anonymous):

ok so now ill have \[\frac{ 9z-15 }{z^2-z-12 }\] right?

OpenStudy (karatechopper):

@jevans17 Please refrain from posting your question inside someone else's question. That is just plain Disrespectful.

OpenStudy (anonymous):

did i combine it right?

OpenStudy (anonymous):

is that my final answer?

OpenStudy (anonymous):

(っ⌒‿⌒)っ is it the final answer?(っ⌒‿⌒)っ

hero (hero):

It is correct. That's why I gave you a medal

OpenStudy (anonymous):

im sad its not an answer could it be

hero (hero):

Yes, that is it bro. Remember: (z - 4)(z + 3) = z^2 - z - 12

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (anonymous):

(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ

OpenStudy (anonymous):

(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ(っ⌒‿⌒)っ